# Conceptualization, Data Curation

### v2

@@ 5.5 Explainability
AI models have become crucial tools for condition monitoring of industrial processes, providing unprecedented advancements in efficiency optimization, prediction performance, and data-driven decision making [103]. However, the integration of AI has not been without drawbacks. A significant concern arises from the inherent complexity of AI models, often treated as black boxes due to their inexplicable internal mechanisms. The inability to decipher the rationale behind AI predictions and decisions has led to a lack of trust, hindering their effective real-world implementation. In industrial contexts, where decisions have tangible consequences, the absence of interpretability becomes a critical barrier [137]. Explainable AI (XAI) methods address the interpretability issues of AI and other black-box models by providing end-users with transparent and understandable information regarding the model’s internal decision mechanisms.
@@
XAI is an area of research focused on clarifying the intricate functionality of AI models and illuminating their decision-making processes. In essence, XAI bridges the gap between complex AI methodologies and human comprehensibility. By examining a model’s outputs and unravelling its internal workings, XAI techniques empower users with insights into the rationale behind AI-generated outcomes. Many XAI techniques have been explored in the literature, though they can generally be classified as either intrinsic approaches, where the model is designed or refined to be particularly interpretable by humans, or post-hoc approaches, where external tools are applied in parallel to analyze the model’s internal mechanisms with respect to predicted outputs [61]. Intrinsic methods, sometimes referred to as interpretable AI, are concerned with the development of the AI model itself, limiting the design to simple methods such as linear regression and logic based rules. Post-hoc methods on the other hand, can be model specific or general to any Blackbox model with varying levels of explanations either locally or globally [18]. Commonly used post-hoc XAI tools include local interpretable model-agnostic explanations (LIME) [76], which works by learning local interpretable models around the prediction, and Shapley additive explanations (SHAP) [50], which uses ideas from game theory to explain model contributions.
@@
In the study by Bhakte et al. [7], an XAI-based condition monitoring framework is introduced specifically for understanding the predictions made by deep learning-based FDD models. The application of this research provides two main benefits: it enhances user confidence by providing explanations for correct fault identifications and it empowers model developers with insights when the model identifies faults incorrectly. Their methodology, based on the Shapley value framework, uses integrated gradients to identify the variables influencing DNN predictions, which are often related to the root causes of faults. A similar study [70] presents a comprehensive two-step framework for the detection and diagnosis of faults in industrial processes, with an emphasis on providing better understandability. The first step leverages a denoising sparse autoencoder to acquire a robust and adaptable feature embedding, while the subsequent phase employs smooth integrated gradients to unveil the significance of features tied to detected faults. Additionally, [2] presents an innovative deep learning-based FDD method that harnesses XAI to refine input features based on their relevance. By utilizing the layerwise relevance propagation (LRP) algorithm, this approach not only eliminates irrelevant input variables but also offers potential insights into the root causes of detected faults. Collectively, these studies underscore the profound impact of XAI in enhancing the transparency, dependability, and interpretability of condition monitoring models, ultimately yielding more effective fault detection, diagnosis, and comprehension.
5.6 Causality
@@
Causality plays a key role in the condition monitoring of industrial processes, offering a better understanding of the true cause-and-effect relationships within complex systems. While XAI offers insights into the reasoning and inherent behaviour of intelligent models and their outputs, it neglects to consider the true physical systems that generate the data. In contrast, causality is an area of research that specifically focuses on understanding the true causal mechanisms driving physical systems rather than simply relying on statistical associations between variables [64]. These underlying relationships hold particular significance for industrial processes. Industrial processes are often characterized by a multitude of process variables and parameters, coupled with external disturbances and noise. Consequently, data-driven methods are highly susceptible to spurious correlations and erroneous associations. Thus, the exploration of causal analysis becomes critical for analyzing intricate industrial processes.
@@
Causal modelling is the formal representation of these cause-and-effect relationships, often expressed through Bayesian networks or structural equation models. These models are generally crafted by domain experts using their expertise and leveraging their intimate knowledge of the systems. However, when domain knowledge is unavailable, causal models can be approximated via randomized control trials and experimental interventions. Even so, analyzing industrial processes in these ways might prove unfeasible or unsafe. To circumvent this problem, causal discovery methods, such as those examined by Vukovic et al. [88], become indispensable. These methods employ statistical algorithms to analyze historical observational data and infer causal relationships among the complex web of variables.
@@
The integration of causality into data-driven AI referred to as Causal AI (CAI), can provide more robust and interpretable model predictions. Since causality represents the fundamental data generation mechanisms of a system, CAI models are less biased toward training data and are more effective in new situations by applying causal generalizations from past observations [19]. Causality-informed AI methods can provide more actionable insights by pinpointing the actual causes of anomalies or issues, facilitating targeted and effective responses. Additionally, CAI approaches can reduce false positives/negatives by distinguishing between spurious correlations and true causal links. Furthermore, these methods can be easily adapted for root-cause analysis, an important aspect of fault diagnosis and condition monitoring.
@@
Several notable studies demonstrate the practical application of CAI in industrial process condition monitoring. For instance, Ma et al. [52] have developed a KPI-oriented hierarchical monitoring framework that utilizes Bayesian fusion and a Granger causality method to enhance fault detection and illuminate propagation paths. Similarly, Wu et al. [95] propose a multimodal monitoring framework for fault detection and propagation analysis using a deep local adaptive network (DLAN) alongside a dedicated causal Bayesian network. Li et al. [42] employ a spatial sequence-based deep learning approach, leveraging LSTM models with attention mechanisms to predict faults and identify causal variables with remarkable accuracy. Arunthavanathan et al. [5] present an autonomous framework for online monitoring that continuously updates itself, integrating causality through variable contributions extracted by a permutation algorithm. Meanwhile, Bi et al. [9] introduce a novel causal discovery method centred on the causality-gated time-series transformer (CGTST) model, explicitly identifying causal variables and refining relationships through permutation feature importance (PFI). In contrast, Wang et al. [91] harness causality with their temporal registration network (TRN), demonstrating improved root cause localization and propagation path identification. These select studies demonstrate the potential of integrating causality to advance the field of industrial process condition monitoring, offering comprehensive insights for more effective fault detection and diagnosis.
@@ 6 Conclusions
This paper offers an insightful examination of the utilization of deep learning and machine learning algorithms for industrial plant condition monitoring. Its primary focus centres on deep learning-based methods, renowned for their exceptional accuracy and resilience in this context. The Tennessee Eastman Process serves as the principal benchmark for evaluation, and the conclusions that follow apply to similar facilities in the chemical industry as well as complex systems that are facing inherent uncertainty. This paper emphasises how crucial deep learning algorithms are to making the complex process of feature extraction simpler. These algorithms may automatically identify and extract critical features without the need for manual feature engineering or prior experience by utilising state-of-the-art methods such as autoencoders. Furthermore, this study recognises and tackles the significant difficulties associated with managing unlabeled and imbalanced data, which are widespread problems in actual industrial monitoring situations. In addition to identifying these difficulties, it offers insightful information on how to navigate and overcome them, strengthening the usefulness and resilience of deep learning-based condition monitoring in industrial settings. To enhance the interpretability of the AI systems and make the decisions made by ML algorithms more trusted, Explainable AI and Causal AI methods are explored. In essence, this research advances industrial condition monitoring by clarifying the benefits of deep learning methods, highlighting their ability to simplify complicated problems, and providing workable answers to industry problems. The results showcased here facilitate industrial plant condition monitoring, benefiting both researchers and practitioners.

### v3

@@ 4.2.3 Explainability
AI models have become crucial tools for condition monitoring of industrial processes, providing unprecedented advancements in efficiency optimization, prediction performance, and data-driven decision making Yan et al. (2023). However, the integration of AI has not been without drawbacks. A significant concern arises from the inherent complexity of AI models, often treated as black boxes due to their inexplicable internal mechanisms. The inability to decipher the rationale behind AI predictions and decisions has led to a lack of trust, hindering their effective real-world implementation. In industrial contexts, where decisions have tangible consequences, the absence of interpretability becomes a critical barrier Zio (2022). Explainable AI (XAI) methods address the interpretability issues of AI and other black-box models by providing end-users with transparent and understandable information regarding the model’s internal decision mechanisms.
@@
XAI is an area of research focused on clarifying the intricate functionality of AI models and illuminating their decision-making processes. In essence, XAI bridges the gap between complex AI methodologies and human comprehensibility. By examining a model’s outputs and unraveling its internal workings, XAI techniques empower users with insights into the rationale behind AI-generated outcomes. Many XAI techniques have been explored in the literature. However, they can generally be classified as either intrinsic approaches, where the model is designed or refined to be particularly interpretable by humans, or post-hoc approaches, where external tools are applied in parallel to analyze the model’s internal mechanisms with respect to predicted outputs Nor et al. (2021). Intrinsic methods, sometimes referred to as interpretable AI, are concerned with the development of the AI model itself, limiting the design to simple methods such as linear regression and logic-based rules. Post-hoc methods, on the other hand, can be model-specific or general to any Black box model with varying levels of explanations, either locally or globally Confalonieri et al. (2021). Commonly used post-hoc XAI tools include local interpretable model-agnostic explanations (LIME) Ribeiro et al. (2016), which works by learning local interpretable models around the prediction, and Shapley additive explanations (SHAP) Lundberg and Lee (2017), which uses ideas from game theory to explain model contributions.
@@
In the study by Bhakte et al. Bhakte et al. (2022), an XAI-based condition monitoring framework is introduced specifically for understanding the predictions made by deep learning-based FDD models. The application of this research provides two main benefits: it enhances user confidence by providing explanations for correct fault identifications, and it empowers model developers with insights when the model identifies faults incorrectly. Their methodology, based on the Shapley value framework, uses integrated gradients to identify the variables influencing DNN predictions, which are often related to the root causes of faults. A similar study Peng et al. (2022d) presents a comprehensive two-step framework for the detection and diagnosis of faults in industrial processes, with an emphasis on providing better understandability. The first step leverages a denoising sparse autoencoder to acquire a robust and adaptable feature embedding, while the subsequent phase employs smooth integrated gradients to unveil the significance of features tied to detected faults. Additionally, Agarwal et al. (2021) presents an innovative deep learning-based FDD method that harnesses XAI to refine input features based on their relevance. By utilizing the layerwise relevance propagation (LRP) algorithm, this approach not only eliminates irrelevant input variables but also offers potential insights into the root causes of detected faults. Rule-based classification methods like Logical Analysis of Data (LAD) are used for interpretable fault detection in complex industrial processes like TEP Ragab et al. (2018), where hidden knowledge is discovered by revealing interpretable patterns linked to underlying physical phenomena. The decision model explains the potential causes of these faults. Another method to increase the interpretability and explainability of CM, especially for complex systems, is using physics-informed machine learning (PIML) approaches Wu et al. (2024) that combine physics-based modeling and data-driven machine learning. Collectively, these studies underscore the profound impact of XAI in enhancing the transparency, dependability, and interpretability of condition monitoring models, ultimately yielding more effective fault detection, diagnosis, and comprehension.
To enhance the explainability of Graph Convolutional Network (GCN)-based fault diagnosis methods in industrial systems, Graph-SRGA, a two-stage variable contribution analysis framework, is proposed in Xu et al. (2025). Traditional GCN-based approaches treat fault diagnosis as a pure classification task, often neglecting the interpretability of which variables or sensor nodes most influence diagnostic outcomes. This restricts industrial deployment due to reliability concerns. Graph-SRGA introduces a two-stage analysis process: Neighbor-stage contribution analysis assesses the impact of neighboring nodes in the graph. Sensor-stage contribution analysis evaluates the role of individual sensors connected to these nodes. The method employs a self-reproductive non-dominated sorting genetic algorithm (SRGA), guided by two objective functions, to identify influential variable subsets that contribute most significantly to the GCN’s decision-making. Experimental validation on the TEP and high-speed train traction systems confirms that Graph-SRGA provides accurate, interpretable insights into the contributions of various features, offering a practical solution for improving the transparency and trustworthiness of GCN-based diagnostic models in industrial applications.
4.2.4 Causality
@@
Causality plays a key role in the condition monitoring of industrial processes, offering a better understanding of the true cause-and-effect relationships within complex systems. While XAI provides insights into the reasoning and inherent behavior of intelligent models and their outputs, it neglects to consider the true physical systems that generate the data. In contrast, causality is an area of research that specifically focuses on understanding the true causal mechanisms driving physical systems rather than simply relying on statistical associations between variables Pearl (2009). These underlying relationships hold particular significance for industrial processes. Industrial processes are often characterized by a multitude of process variables and parameters, coupled with external disturbances and noise. Consequently, data-driven methods are highly susceptible to spurious correlations and erroneous associations. Thus, the exploration of causal analysis becomes critical for analyzing intricate industrial processes.
@@
Causal modeling is the formal representation of these cause-and-effect relationships, often expressed through Bayesian networks or structural equation models. These models are generally crafted by domain experts using their expertise and leveraging their intimate knowledge of the systems. However, when domain knowledge is unavailable, causal models can be approximated via randomized control trials and experimental interventions. Even so, analyzing industrial processes in these ways might prove unfeasible or unsafe. To circumvent this problem, causal discovery methods, such as those examined by Vukovic et al. Vuković and Thalmann (2022), become indispensable. These methods employ statistical algorithms to analyze historical observational data and infer causal relationships among the complex web of variables.
@@
The integration of causality into data-driven AI referred to as Causal AI (CAI), can provide more robust and interpretable model predictions. Since causality represents the fundamental data generation mechanisms of a system, CAI models are less biased toward training data and are more effective in new situations by applying causal generalizations from past observations Cox Jr (2023). Causality-informed AI methods can provide more actionable insights by pinpointing the actual causes of anomalies or issues, facilitating targeted and effective responses. Additionally, CAI approaches can reduce false positives/negatives by distinguishing between spurious correlations and true causal links. Furthermore, these methods can be easily adapted for root-cause analysis, an important aspect of fault diagnosis and condition monitoring.
@@
Several notable studies demonstrate the practical application of CAI in industrial process condition monitoring. For instance, Ma et al. Ma et al. (2020) have developed a KPI-oriented hierarchical monitoring framework that utilizes Bayesian fusion and a Granger causality method to enhance fault detection and illuminate propagation paths. Similarly, Wu et al. Wu et al. (2023) propose a multimodal monitoring framework for fault detection and propagation analysis using a deep local adaptive network (DLAN) alongside a dedicated causal Bayesian network. Li et al. Li (2021) employ a spatial sequence-based deep learning approach, leveraging LSTM models with attention mechanisms to predict faults and identify causal variables with remarkable accuracy. Arunthavanathan et al. Arunthavanathan et al. (2022) present an autonomous framework for online monitoring that continuously updates itself, integrating causality through variable contributions extracted by a permutation algorithm. Meanwhile, Bi et al. Bi et al. (2023) introduce a novel causal discovery method centred on the causality-gated time-series transformer (CGTST) model, explicitly identifying causal variables and refining relationships through permutation feature importance (PFI). In contrast, Wang et al. Wang et al. (2023b) harness causality with their temporal registration network (TRN), demonstrating improved root cause localization and propagation path identification. These select studies demonstrate the potential of integrating causality to advance the field of industrial process condition monitoring, offering comprehensive insights for more effective fault detection and diagnosis.
@@
4.3 Industrial Applicability Beyond TEP
While the TEP benchmark provides a standardized evaluation, its synthetic nature cannot fully replicate noise patterns, unmodeled dynamics, and operational variability of actual plants. However, data-driven methods have demonstrated strong generalizability, with models developed for one case study often transferable to others with minimal adaptation while maintaining high performance. For example, MRJDCNN-LSTM, introduced in Chen et al. (2024), demonstrated a high accuracy of 97.47% when tested on the F101/3 coking furnace. Similarly, the GLACNN Yang et al. (2025) validated on penicillin fermentation processes, achieving an FDR of 97.5%. The 1-DCNN + SDAE introduced in Zhang et al. (2021) was validated by the fed-batch fermentation penicillin process (FBFP), and a real-life manufacturing case of an industrial conveyor belt achieving accuracies of 99.68% and 95.0%, respectively. The effectiveness of the unsupervised data mining method introduced in Zheng and Zhao (2020) was also evaluated by an industrial hydrocracking process. The unsupervised ensemble of autoencoders Plakias and Boutalis (2022) was tested on the Case Western Reserve University (CWRU), the Intelligent Maintenance System (IMS) bearing datasets alongside the TEP. BRNN Sun et al. (2020) was validated using a real chemical manufacturing dataset, specifically an amine tower, demonstrating its superiority in extracting data patterns from complex real-world chemical plants.
@@ 5 Conclusions
This paper provides a comprehensive review of the application of machine learning (ML) and deep learning (DL) techniques for condition monitoring in industrial plants, with a particular emphasis on chemical processes and the Tennessee Eastman Process (TEP) benchmark. The survey highlights the growing dominance of DL-based methods, which have demonstrated superior capability in handling complex, high-dimensional process data and in automating feature extraction—reducing reliance on manual engineering and domain-specific expertise. Key challenges frequently encountered in real-world industrial monitoring, including imbalanced and unlabeled data, are critically discussed alongside strategies such as advanced data augmentation, semi-supervised learning, and ensemble methods. The study also highlights the importance of explainability and causal reasoning in AI-driven monitoring systems, as these elements are essential for enabling trust, interpretability, and adoption in safety-critical domains. From a practical standpoint, this work emphasizes that the effective deployment of AI-based condition monitoring has direct implications for process safety, risk management, and operational efficiency. By outlining emerging solutions and research directions, this review not only consolidates current knowledge but also identifies pathways toward robust, transparent, and industrially viable monitoring frameworks. In conclusion, intelligent condition monitoring supported by DL and ML represents a transformative approach for industrial systems, offering improved fault detection, reduced downtime, enhanced safety, and greater reliability. Continued efforts in addressing uncertainty, improving explainability, and validating hybrid and transferable models across diverse plants will be pivotal for advancing the next generation of industrial monitoring solutions.